OpenStudy (anonymous):

Two liters of a perfect gas are at 0 degree C and 1 atm.If the gas is nitrogen, N2. Determine the mass of the gas?

OpenStudy (anonymous):

Use the ideal gas law\[PV = nRT\] to determine the amount of the gas, n. Then we know that\[{\rm mass }= n \cdot MM\] where MM is the molar mass. Use the following value for R. \[R = 0.08205746 \left [ \rm L \cdot atm \over K \cdot mol \right ]\] Remember to add 273 to the temperature value in Celsius to get it into Kelvin.